# Adler function, Bjorken polarized Sum Rule – confirmation of elements of the $\{\beta\}$-expansion

### v1

@@ 1 Introduction
Here we will consider the results of $\{\beta\}$–expansion Mikhailov:2004iq and discuss the reliability of the obtained elements of this expansion Baikov:2022zvq and the relation of the latter with the corresponding Feynman diagrams. The elements of this expansion appear at the decomposition of the standard perturbative coefficients of any renormalization group invariants (RGIs); here these coefficients are $d_{n}$ or $c_{n}$ for the nonsinglet Adler $D_{A}$-function, or the Bjorken polarized sum rule (BpSR) $S^{Bjp}$ respectively. We consider independent confirmation of our results for the decomposition of the nonsinglet Adler $D_{A}$-function for the values of elements $d_{3}[.]$ ($c_{3}[.]$) of $\{\beta\}$–expansion and for some elements of $d_{4}[.]$ Baikov:2022zvq . The agreements are obtained using the already known results of N3,4LO calculations Ball:1995ni of some diagrams in the $\overline{\rm MS}$scheme in pure QCD. In the case of the Adler nonsinglet $D_{A}$-function at the default choice of scale $Q^{2}=\mu^{2}$,
@@
where $\beta_{i}$ are the coefficients of expansion of the QCD $\beta$-function (see Appendix A). The $\{\beta\}$-expansion shows how the intrinsic charge renormalization manifests itself; thus all possible $\beta_{i}$-terms appear in each order $n$ of perturbative expansion (PT). This property differ it from more or less formal expansions over powers of selected lagrangian parameters or Casimirs, say the number of active quark flavors $n_{f}$, or powers of $\beta_{0}$ Lovett-Turner:1995zwc , which is based on the same formal reasons. The procedure of calculating the expansion elements in Eq. (2) was based, in its turn, on the results of extended QCD (QCDe) presented in Chetyrkin:2022fqk , which possesses a number of new degrees of freedom (d.o.f.) revealing themselves in intrinsic loops. After decomposition as in Eq.(2), one can equate the contributions of these new d.o.f. to zero and return from QCDe to the standard QCD. Note here that now there are two different approaches to obtain values of the elements within the standard QCD only, which lead to different results. One of them, which started with Mojaza:2012mf , is based on an interpretation of RG transformation, it got the name “Principle of Maximum Conformality” (PMC), see review DiGiustino:2023jiq and references therein. The other approach, see Cvetic:2016rot , Goriachuk:2021ayq , is based on an extended interpretation of the Crewter-Broadharst-Kataev (CBK) relation Broadhurst:1993ru . So some doubts appear about how reliable the results for the elements obtained within QCDe in Baikov:2022zvq are with respect to the standard QCD.
Here we confirm our results in order $O(a_{s}^{3})$ and confirm them in part in order of $O(a_{s}^{4})$ based on the calculation within the standard QCD and focusing on the diagrammatic origins of the elements of the $\{\beta\}$-expansion. In addition in Sec.4 we propose a hierarchy of contributions to the coefficients $d_{n}(c_{n})$ based on a certain generalization of Naive Nonabelianization (NNA) in the framework of the $\{\beta\}$-expansion. The usage of this hierarchy can facilitate the laborious process of estimating physical quantities in high-lows calculations.
@@ 2 Test for $\{\beta\}$-expansion elements of $D_{A}\,(S^{Bjp})$ at $O(a_{s}^{3})$
These decompositions for $d_{3}$ were first obtained in Mikhailov:2004iq where one used the results for $D_{A}$ with a single additional degree of freedom – the light gluino Chetyrkin:1996ez , then the elements of this decomposition were confirmed in Baikov:2022zvq within QCDe with any numbers of d.o.f. Below we reproduce the same values of the elements $d_{3}[.]$ by considering the results for specific diagrams for $D_{A}$ calculated in pure QCD.
Let us start with the simplest case in (3). The first element $d_{2}[1]$ originates from the diagram where the single gluon line is dressed in a quark bubble, wich is proportional to $T_{R}n_{f}\sim\beta_{0}$. Just this correspondence allows one to perform the decomposition unambiguously. The sequential dressing of this only one gluon line generates its anomalous dimension that is finally revealed as a set of $\beta$-elements of our interest also in the higher orders of PT.
@@
We reproduced in Fig.1 the diagrams taken from Fig.C1(IV1) and Fig.C1(IV2) in Ball:1995ni , and the abelian projection of the results corresponding to them was presented in Appendix C there. The result for the diagram in Fig.1(Left) independently confirms the value of the $d_{3}[2]$ term in (4a) and coincides also with the general expression for $d_{n}[n-1]$, which can be extracted from the results of Broadhurst:1992si ; Broadhurst:1993ru . The diagram in the right panel contributes to the term $\beta_{1}d_{3}[0,1]$ by means of anomalous dimension of a gluon in order of $a_{s}^{2}$. The latter, in its turn, includes the unique Casimir $C_{\text{F}}\cdot(T_{R}n_{f}\equiv x)$, which is common with $\beta_{1}$, see Eq.(21b). This diagrammatic correspondence unambiguously determines the element $d_{3}[0,1]$ within the standard QCD calculation. Further we will trace the contributions to $d_{n}$ of the color structures starting with maximum numbers of quark loops: the evident case of a renormalon chain with only one gluon line contribution gives us $C_{F}\cdot x^{(n-1)}$, then one fermion loop less – $C_{F}(C_{F}x^{(n-2)})$, then $C_{F}(C_{F}^{2}x^{(n-3)})$ and so on. In other words, these terms are a projection onto the abelian sector of the QCD calculation (proportional to QED) of $D$ and $S^{Bjp}$. The abelian part of the structure of $\beta$-coefficients is well recognized in the first terms (underlined) of the corresponding formulas in Appendix A.
@@ 3 Test for $\{\beta\}$-expansion elements of $D_{A}$ at $O(a_{s}^{4})$
one should supplement $s_{4}$ with the contribution $\hat{P}_{C^{2}_{F}x^{2}}\left(\beta_{0}^{2}d_{4}[2]\right)$ in the rhs. The diagrams forming this additional term are not of “the only one dressed gluon line” type. Equality (10) unambiguously determines the abelian part of the element $d_{4}[2]$ and confirms now Eq.(8d). Concluding we confirm the values of the elements $d_{4}[0,0,1],d_{4}[1,1]$ and the abelian part of $d_{4}[2]$ based on the direct QCD calculations  [^2] The corresponding values of $d_{4}[0,0,1],d_{4}[1,1]$ presented in Cvetic:2016rot , and Kataev:2022iqf , Goriachuk:2021ayq (Table 2 there) do not agree with the lhs in (3), and the abelian part of $d_{4}[2]$ does not agree with (10).
@@
Let us mention that if the elements $d_{4}[0,0,1],d_{4}[0,1]$ are already known, then the counter part elements $c_{4}[0,0,1],c_{4}[0,1]$ are also known from the corollaries Baikov:2022zvq of the CBK relation Baikov:2022zvq ; Chetyrkin:2022fqk like Eq.(7).
@@ 4.1 Generalized NNA, illustration of $\beta_{0}$-dominance
Then, we follow the “practical” observation $\beta_{i}=O(\beta_{0}^{i+1})$ that works at $n_{f}=0\div 5$, $i=1\div 4$, and invent the notation $b_{i}\stackrel{{\scriptstyle def}}{{=}}\beta_{i}/\beta_{0}^{i+1}=O(1)$, see (23). Of course, the last equality should be broken down at higher orders $i$ of pQCD due to the factorial growth of $\beta_{i}$. So joining both these conditions, we take into account the terms with the coefficients $b_{i}$ together with others in the same order of $1/(\beta_{0})^{i}$, see the first brackets in Eqs.(14) and the first and the second brackets in Eqs.(15) below. We call this kind of rearrangement in the groups of the same $1/\beta_{0}$-order in brackets the generalized NNA (gNNA). The expansion coefficients of the $D_{A}$-function and BpSR $S^{Bjp}$ are presented below as the sums of these groups accompanied by their numerical estimates (at $n_{f}=3$, NNA estimates are shown from below in red),
$d_{2}=\beta_{0}\Big{[}\underbrace{d_{2}[1]}_{{\color[rgb]{1,0,0}\bm{2.77}}}+\overbrace{\frac{1}{\beta_{0}}d_{2}[0]}^{\bm{0.15}}\Big{]}\approx\beta_{0}[\bm{2.92}];~{}c_{2}=\beta_{0}\Big{[}\underbrace{c_{2}[1]}_{{\color[rgb]{1,0,0}\bm{-8}}}+\overbrace{\frac{1}{\beta_{0}}d_{2}[0]}^{\bm{1.6}}\Big{]}\approx\beta_{0}[\bm{-6.4}]$ (13)
@@
As it is seen, the leading $(1/\beta_{0})^{0}$-groups numerically dominate in all of the sums in (14,15), while the subleading groups for the most part mutually cancel one another, compare the lhs and the rhs there. The evident manifistation of gNNA vs NNA can be seen in the leading group of (15a). The numerical estimates of each of the elements $d_{3,4}[.],c_{3,4}[.]$ are shown in Appendix B, where the diagrammatically confirmed results are marked in red. In addition, just the result of this leading group has been independently confirmed here in Sec.3, see the rhs of Eq.(3), which was based on the calculations of the QCD diagrams presented in Fig.2. Therefore, the correspondence of gNNA and the total sum in (15a) does not even depend on the detailed results in Eqs.(8b,8c) obtained with the help of QCDe. On the other hand, the contributions of the suppressed in $1/\beta_{0}$ the so-called “scale invariant” terms $d_{3,4}[0]$ almost cancel with the other subleading groups. At the same time, these $d_{i}[0](c_{i}[0])$ do not correspond by themselves to the final sum in the rhs neither in value nor in sign  [^4] The authors of PMC, see DiGiustino:2023jiq and references therein, propagate another, important role of the “scale invariant” elements $d_{n}[0]$ within the PMC approach.. We recognize that our conjecture is based on observations of a few examples only, and the reason of the aforementioned mutual cancellations is not clear; nevertheless, it looks reasonable to formulate the conjecture for the next orders of pQCD.
@@ 4.2 Conjecture of $\beta_{0}$-dominance
The number of elements of the leading group in Eq.(16c) is equal to the number of partitions $p$ depending on perturbation order $n$, e.g. , here – $p(5-1)=\bm{5}$ OEIS , see the detailed discussion of this correspondence in Baikov:2022zvq , Sec.4. So taking a natural series for the argument of the function $p$, one obtains $p(1,2,3,\bm{4},5,6,7,\ldots)=$ $1,2,3,\bm{5},7,11,15,\ldots$ For the order $n$ of pQCD the leading gNNA group consists of the number of partitions $p(n-1)$ elements that can be traced from the dressing of only one gluon line of the diagram. The indices of the elements $d_{n}[j_{1},\ldots,j_{n-1}]$ of the leading group satisfy the evident equality $11j_{1}+2j_{2}+\ldots\!+\!(\!n\!-1)\!j_{n-1}=n-$ 1, while for the different subleading groups the conditions for the indices are $1\,j_{1}+2\,j_{2}+\ldots+(n-1)\,j_{n-1}=n-2,n-3,\ldots$, where the shift from the number $n-1$ in the rhs determines the order of $1/\beta_{0}$ suppression see, e.g. , the groups in (16c, 16c).
@@
sum of diag.(b,c,d) $=$ $P_{C_{\text{F}}x^{3}}\left(\beta_{3}d_{5}[0,0,0,1]+\beta_{2}\beta_{0}\,d_{5}[1,0,1]+\beta_{1}\beta_{0}^{2}\,d_{5}[2,1]\right)$ (17)
sum of diag.(e,f) $=$ $P_{C_{\text{F}}^{2}x^{2}}\left(\beta_{3}d_{5}[0,0,0,1]+\beta_{1}^{2}\,d_{5}[0,2]\right)$ (18)
@@ 5 Conclusion
We have confirm our results for the elements of the $\{\beta\}$-expansion in Baikov:2022zvq ; Mikhailov:2004iq : in order of $O(a_{s}^{3})$ – completely; in order of $O(a_{s}^{4})$ we confirm two elements, $d_{4}[0,0,1],d_{4}[1,1]$, and the abelian part of $d_{4}[2]$ from Baikov:2022zvq , and for both the cases – basing on the known calculation Ball:1995ni within the standard QCD only.

### v3

# Adler function, Bjorken polarized Sum Rule: confirmation of elements of the $\{\beta\}$-expansion

@@ 1 Introduction
Here we will consider the results of $\{\beta\}$–expansion Mikhailov:2004iq and discuss the reliability of the obtained elements of this expansion Baikov:2022zvq and the relation of the latter with the corresponding Feynman diagrams. The elements of this expansion appear at the decomposition of the standard perturbative coefficients of any renormalization group invariants (RGIs); here these coefficients are $d_{n}$ or $c_{n}$ for the nonsinglet Adler $D_{A}$-function, or the Bjorken polarized sum rule (BpSR) $S^{Bjp}$ respectively. We consider independent confirmation of our results for the decomposition of the nonsinglet Adler $D_{A}$-function for the values of elements $d_{3}[.]$ ($c_{3}[.]$) of $\{\beta\}$–expansion and for some elements of $d_{4}[.]$ Baikov:2022zvq . The agreements are obtained using the already known results of N3,4LO calculations Ball:1995ni of some diagrams in the $\overline{\rm MS}$scheme in pure QCD. In the case of the nonsinglet $D_{A}$-function at the default choice of scale $Q^{2}=\mu^{2}$,
@@
where $\beta_{i}$ are the coefficients of expansion of the QCD $\beta$-function (see Appendix A). The $\{\beta\}$-expansion shows how the intrinsic charge renormalization manifests itself; thus all possible $\beta_{i}$-terms appear in each order $n$ of perturbative expansion (PT). This property differ it from more or less formal expansions over powers of selected lagrangian parameters or Casimirs, say the number of active quark flavors $n_{f}$, or powers of $\beta_{0}$ Lovett-Turner:1995zwc , which is based on the same formal reasons. The procedure of calculating the expansion elements in (2) was based, in its turn, on the results of extended QCD (QCDe) presented in Chetyrkin:2022fqk , which possesses a number of new degrees of freedom (d.o.f.). These d.o.f. are arbitrary number of different fermion representations, revealing themselves only in intrinsic loops. After decomposition as in (2), one can equate the contributions of these new d.o.f. to zero and return from QCDe to the standard QCD. Note here that now are at least two different approaches to obtain the values of the elements within the standard QCD that lead to different results. One of them, started with Mojaza:2012mf , is based on an interpretation of RG transformation, it got the name “Principle of Maximum Conformality” (PMC), see review DiGiustino:2023jiq and references therein. The other approach, see Cvetic:2016rot , is based on an extended interpretation of Crewther-Broadhurst-Kataev (CBK) relation Broadhurst:1993ru . So, some doubt appeared about how reliable the results for the elements obtained within QCDe in Baikov:2022zvq are with respect to the standard QCD.
Here we confirm our results in order $O(a_{s}^{3})$ and confirm them in part in order of $O(a_{s}^{4})$ based on the calculation within the standard QCD and focusing on the diagrammatic origins of the elements of the $\{\beta\}$-expansion. In addition in Sec.4 we propose a hierarchy of contributions to the coefficients $d_{n}(c_{n})$ based on a certain generalization of Naive Nonabelianization (NNA) in the framework of the $\{\beta\}$-expansion. The usage of this hierarchy can facilitate the laborious process of estimating physical quantities in high-loops calculations.
@@ 2 Test for $\{\beta\}$-expansion elements of $D_{A}\,(S^{Bjp})$ at $O(a_{s}^{3})$
These decompositions for $d_{3}$ were first obtained in Mikhailov:2004iq where one used the results for $D_{A}$ with a single additional degree of freedom – the light gluino Chetyrkin:1996ez , then the elements of this decomposition were confirmed in Baikov:2022zvq within QCDe with any numbers of fermion d.o.f. Below we reproduce the same values of the elements $d_{3}[.]$ by considering the results for specific diagrams for $D_{A}$ calculated in pure QCD.
Let us start with the simplest case in (3). The first element $d_{2}[1]$ originates from the diagram where the single gluon line is dressed in a quark bubble, which is proportional to $T_{R}n_{f}\sim\beta_{0}$. Just this correspondence allows one to perform the decomposition unambiguously. The sequential dressing of this only one gluon line generates its anomalous dimension that is finally revealed as a set of $\beta$-elements of our interest also in the higher orders of PT.
@@
We reproduced in Fig.1 the diagrams taken from Fig.C1(IV1) and Fig.C1(IV2) in Ball:1995ni , and the abelian projection of the results corresponding to them was presented in Appendix C, Eqs.(C.2),(C.3) there. The result for the diagram in Fig.1(Left) independently confirms the value of the $d_{3}[2]$ term in (4a) and coincides also with the general expression for $d_{n}[n-1]$, which can be extracted from the results of Broadhurst:1992si ; Broadhurst:1993ru . The diagram in the right panel contributes to the term $\beta_{1}d_{3}[0,1]$ by means of anomalous dimension of a gluon in order of $a_{s}^{2}$. The latter, in its turn, includes the unique Casimir $C_{\text{F}}\cdot(T_{R}n_{f}\equiv x)$, which is common with $\beta_{1}$, see Eq.(24b). This diagrammatic correspondence unambiguously determines the element $d_{3}[0,1]$ within the standard QCD calculation that value coincides with the one in (4a). Further we will trace the contributions to $d_{n}$ of the color structures starting with maximum numbers of quark loops: the evident case of a renormalon chain with only one gluon line contribution gives us $C_{F}\cdot x^{(n-1)}$, then one fermion loop less – $C_{F}(C_{F}x^{(n-2)})$, then $C_{F}(C_{F}^{2}x^{(n-3)})$ and so on. In other words, these terms are a projection onto the abelian sector of the QCD calculation (proportional to QED) of $D_{A}$ and $S^{Bjp}$. The abelian part of the structure of $\beta$-coefficients is well recognized in the first terms (underlined) of the corresponding formulas in Appendix A.
@@ 3 Test for $\{\beta\}$-expansion elements of $D_{A}$ at $O(a_{s}^{4})$
one should supplement $s_{4}$ with the contribution $\hat{P}_{C^{2}_{F}x^{2}}\left(\beta_{0}^{2}d_{4}[2]\right)$ in the rhs. The diagrams forming this additional term are not of “the only one dressed gluon line” type. Equality (10) unambiguously determines the abelian part of the element $d_{4}[2]$ and confirms now Eq.(8d). Concluding we confirm the values of the elements $d_{4}[0,0,1],d_{4}[1,1]$ and the abelian part of $d_{4}[2]$ based on the direct QCD calculations  [^2] The corresponding values of $d_{4}[0,0,1],d_{4}[1,1]$ presented in Cvetic:2016rot , and Kataev:2022iqf , Goriachuk:2021ayq (Table 2 there) do not agree with the lhs in (3), the abelian part of their $d_{4}[2]$ does not agree with (10) with this $s_{4}$..
@@
It is important if the elements $d_{4}[0,0,1],d_{4}[0,1]$ are already known, then the counter part elements $c_{4}[0,0,1],c_{4}[0,1]$ are also known from the corollaries Baikov:2022zvq of the CBK relation Baikov:2022zvq ; Chetyrkin:2022fqk like Eq.(7).
Let us mention here a rather rough test for elements $d_{3,4}[0]$ that is related to a special choice of Casimirs $C_{A},C_{F}$, see (25). The coefficients of perturbative expansion based on $SU_{c}(N)$ dynamics are the polynomials over the $C_{A},C_{F}$, let us consider these Casimirs as variables, which are related by the condition $07C_{A}+11C_{F}=$ 0 (that formally corresponds to the value $N=\sqrt{11}/5$). This linear condition, see Eqs.(5b, 5d), leads to the evident nullification of the coefficients $\beta_{0},\beta_{1}$,
$\beta_{0}(x_{0})=\beta_{1}(x_{0})=\beta_{0}(x_{1})=\beta_{1}(x_{1})=0\,.$ (13)
This condition radically simplifies the color structure of elements and in virtue of (13) leads
$\!\!\!\!\!\text{at}~{}7C_{A}+11C_{F}=0:$
[…]
The Eqs.(4c) for $d_{3}[0]$, and (8c) for $d_{4}[0,0,1]$, (8g) for $d_{4}[0]$ as well as the corresponding results in Goriachuk:2021ayq satisfy both the tests in Eqs.(14,15).
@@ 4.1 Generalized NNA, illustration of $\beta_{0}$-dominance
Then, we follow the “practical” observation $\beta_{i}=O(\beta_{0}^{i+1})$ that works at $n_{f}=0\div 5$, $i=1\div 4$, and invent the notation $b_{i}\stackrel{{\scriptstyle def}}{{=}}\beta_{i}/\beta_{0}^{i+1}=O(1)$, see (27). Of course, the last equality should be broken down at higher orders $i$ of pQCD due to the further factorial growth of $\beta_{i}$. So joining both these conditions, we take into account the terms with the coefficients $b_{i}$ together with others in the same order of $1/(\beta_{0})^{i}$, see the first brackets in Eqs.(17) and the first and the second brackets in Eqs.(18) below. We call this kind of rearrangement in the groups of the same $1/\beta_{0}$-order in brackets the generalized NNA (gNNA). The expansion coefficients of the $D_{A}$-function and BpSR $S^{Bjp}$ are presented below as the sums of these groups accompanied by their numerical estimates in inscriptions (at $n_{f}=3$, NNA estimates are shown from below in red),
$d_{2}=\beta_{0}\Big{[}\underbrace{d_{2}[1]}_{{\color[rgb]{1,0,0}\bm{2.77}}}+\overbrace{\frac{1}{\beta_{0}}d_{2}[0]}^{\bm{0.15}}\Big{]}\approx\beta_{0}[\bm{2.92}];~{}c_{2}=\beta_{0}\Big{[}\underbrace{c_{2}[1]}_{{\color[rgb]{1,0,0}\bm{-8}}}+\overbrace{\frac{1}{\beta_{0}}c_{2}[0]}^{\bm{1.63}}\Big{]}\approx\beta_{0}[\bm{-6.4}]$ (16)
@@
As it is seen, the leading $(1/\beta_{0})^{0}$-groups numerically dominate in all of the sums in (17,18), while the subleading groups for the most part mutually cancel one another, compare the lhs and the rhs there. The evident manifistation of gNNA vs NNA can be seen in the leading group of (18a). The numerical estimates of each of the elements $d_{2,3,4}[.],~{}c_{2,3,4}[.]$ are shown in Appendix B, where the diagrammatically confirmed results are marked in red. In addition, just the result of this leading group has been independently confirmed here in Sec.3, see the rhs of Eq.(3), which was based on the calculations of the QCD diagrams presented in Fig.2. Therefore, the correspondence of gNNA and the total sum in (18a) does not even depend on the detailed results in Eqs.(8b,8c) obtained with the help of QCDe. On the other hand, the contributions of the suppressed in $1/\beta_{0}$ the so-called “scale invariant” terms $d_{3,4}[0]$ almost cancel with the other subleading groups. At the same time, these $d_{i}[0]~{}(c_{i}[0])$ do not correspond by themselves to the final sum in the rhs neither in value nor in sign  [^4] The authors of PMC, see DiGiustino:2023jiq and references therein, propagate another, important role of the “scale invariant” elements $d_{n}[0]$ within the PMC approach.. We recognize that our conjecture is based on observations of a few examples only, and the reason of the aforementioned mutual cancellations is not clear; nevertheless, it looks reasonable to formulate the conjecture for the next orders of pQCD.
@@ 4.2 Conjecture of $\beta_{0}$-dominance
The number of elements of the leading group in Eq.(19c) is equal to the number of partitions $p$ depending on perturbation order $n$, e.g. , here – $p(5-1)=\bm{5}$ OEIS , see the detailed discussion of this correspondence in Baikov:2022zvq , Sec.4. So taking a natural series for the argument of the function $p$, one obtains $p(1,2,3,\bm{4},5,6,7,\ldots)=$ $1,2,3,\bm{5},7,11,15,\ldots$ For the order $n$ of pQCD the leading gNNA group consists of the number of partitions $p(n-1)$ elements that can be traced from the dressing of only one gluon line of the diagram. The indices of the elements $d_{n}[j_{1},\ldots,j_{n-1}]$ of the leading group satisfy the evident equality $11j_{1}+2j_{2}+\ldots\!+\!(\!n\!-1)\!j_{n-1}=n-$ 1, while for the different subleading groups the conditions for the indices are $1\,j_{1}+2\,j_{2}+\ldots+(n-1)\,j_{n-1}=n-2,n-3,\ldots$, where the shift from the number $n-1$ in the rhs determines the order of $1/\beta_{0}$ suppression see, e.g. , the groups in (19c, 19c).
@@
sum of diag.(b,c,d) $=$ $P_{C_{\text{F}}x^{3}}\left(\beta_{3}d_{5}[0,0,0,1]+\beta_{2}\beta_{0}\,d_{5}[1,0,1]+\beta_{1}\beta_{0}^{2}\,d_{5}[2,1]\right);$ (20)
sum of diag.(e,f) $=$ $P_{C_{\text{F}}^{2}x^{2}}\left(\beta_{3}d_{5}[0,0,0,1]+\beta_{1}^{2}\,d_{5}[0,2]\right)\,.$ (21)
@@ 5 Conclusion
We have confirmed our results for the elements of the $\{\beta\}$-expansion in Baikov:2022zvq ; Mikhailov:2004iq : in order of $O(a_{s}^{3})$ – completely; in order of $O(a_{s}^{4})$ we confirmed two elements, $d_{4}[0,0,1],d_{4}[1,1]$, and the abelian part of $d_{4}[2]$ from Baikov:2022zvq , and for both the cases – basing on the known calculation Ball:1995ni within the standard QCD only.